OpenStudy (anonymous):

A balloon has a circumference of 22 in. Use the circumference to approximate the surface area of the balloon to the nearest square centimeter.

OpenStudy (anonymous):

Surface area of a sphere (balloon) is S=4*pi*r^2. So all you need is the radius of the balloon. Since circumference of the balloon is given, we can find the radius using: C=2*pi*r: solving for r, we get r=C/(2*pi) = 22/(2*pi) Now put this value for r into the surface are formula and get your answer... :)